OpenStudy (anonymous):

What is the equation of the line that is perpendicular to -x + y = 7 and passes through (-1, -1)?

OpenStudy (anonymous):

-x + y = 2 -x – y = 2 x – y = 2 -x – y = 0 x + y = 0 options

OpenStudy (dmndlife24):

perpendicular lines have opposite reciprocal slopes right? The new equation will have a slope of -1 The new equation will also pass through (-1, -1) y + 1 = -1(x + 1)

OpenStudy (dmndlife24):

Simplify to get y + 1 = -1x - 1 x + y = -2

OpenStudy (anonymous):

Thats not one of the options

OpenStudy (anonymous):

a

OpenStudy (anonymous):

nevermind I got it thank you